OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (anonymous):

@abhi_abhi @DaWizjr @hlpwntd @agent0smith

OpenStudy (anonymous):

i got A right B and C i have problem on

OpenStudy (anonymous):

a. 7^2 - 2^2 = 49 - 4 = 45 b. 9^9 - 2^2 = 81 - 4 = 77 c. 11^2 - 3^2 = 121 - 9 = 112